# Quantum circuit generation via test-time learning with large language models

### v2

@@ I Introduction
Large language models (LLMs) have become remarkably capable of producing structured artifacts, including executable code, formal proofs, molecular strings, and hardware descriptions, where correctness is not a matter of style but of [5, 21, 10, 19, 29, llm6, llm7, llm8, 9]. Yet the central tension remains: these models are, at their core, stochastic sequence generators. This yielded to the theory of naive sampling approaches like the Large Language Monkeys [3] which, left unattended, can be inventive at the price of high computational resources. A promising direction is to move beyond one-shot prompting and instead scale inference-time [14, 17, 1, 8].
At the same time, test-time learning approaches that equip models with persistent working memory, such as dynamic cheatsheets, suggest a complementary lever: not just sampling repeatedly[3], or back-propagating feedback [28], but query with accumulated experience, to guide the model solution search.
@@
In this work, we explore the potential of a test-learning approach, avoiding the need to prepare a dataset for training or fine-tuning the system. By casting the circuit synthesis as a test-time optimization loop in which an LLM proposes circuit edits from a single initially random circuit. In this way, the generation process becomes a trajectory that can accumulate “skills” specific to the task, such as reusable entangling motifs and topology patterns. We represent circuits directly as gate lists, similarly to other recent jobs that focus on pre-trained or fine-tuning approaches [llm6, 29], and restrict available operations to an experimentally motivated discrete subset, avoiding continuous parameters that would inflate the search space. As the validation metric, we opt for the Meyer–Wallach global entanglement measure, an experimentally-friendly and permutation-invariant metric defined in terms of single-qubit reduced purities. This choice deliberately emphasizes global structure rather than local heuristics: the model is rewarded only when the entire register becomes strongly entangled.
@@ III Main Results
Here, we explore the application of GPT5.1 with an episodic memory step and no feedback. We experiment using 25 to 45 gates on 20-qubit circuits. This choice shrinks the search space, but dodges a trick recurrently applied by small models, which is the application of a high number ($\gg$ num gates) of random CNOTs. Ranging from 25 to 45 gates, we can see from Table 1 that the model improves over the initial random circuit, and can obtain up to perfect MW in some situations, as we can see for 35 gates. Nevertheless, the success probability of having an MW measure $Q(\psi)>0.8$ is $\sim 10\%$. This evidence backs up our insights, demonstrating that LLMs can sample optimal circuits beyond repeated sampling approaches, also known as the Large Language Monkeys [brown2024largelanguagemonkeysscaling]. Nevertheless, to achieve greater performances, a leap is required.
@@ IV Conclusions
Original code used for the experiment and data generated during the experiments are available at: https://github.com/AdriQD/GPT_circuits

### v3

@@ I Introduction
Large language models (LLMs) have become remarkably capable of producing structured artifacts, including executable code, formal proofs, molecular strings, and hardware descriptions, where correctness is not a matter of style but of [4, 23, 9, 20, 32, 12, 19, 28, 8]. Yet the central tension remains: these models are, at their core, stochastic sequence generators. This yielded to the theory of naive sampling approaches like the Large Language Monkeys [brown2024monkey] which, left unattended, can be inventive at the price of high computational resources. A promising direction is to move beyond one-shot prompting and instead scale inference-time [14, 17, 21, 7].
At the same time, test-time learning approaches that equip models with persistent working memory, such as dynamic cheatsheets, suggest a complementary lever: not just sampling repeatedly[brown2024monkey], or back-propagating feedback [31], but query with accumulated experience, to guide the model solution search.
@@
In this work, we explore the potential of a test-learning approach, avoiding the need to prepare a dataset for training or fine-tuning the system. By casting the circuit synthesis as a test-time optimization loop in which an LLM proposes circuit edits from a single initially random circuit. In this way, the generation process becomes a trajectory that can accumulate “skills” specific to the task, such as reusable entangling motifs and topology patterns. We represent circuits directly as gate lists, similarly to other recent jobs that focus on pre-trained or fine-tuning approaches [12, 32], and restrict available operations to an experimentally motivated discrete subset, avoiding continuous parameters that would inflate the search space. As the validation metric, we opt for the Meyer–Wallach global entanglement measure, an experimentally-friendly and permutation-invariant metric defined in terms of single-qubit reduced purities. This choice deliberately emphasizes global structure rather than local heuristics: the model is rewarded only when the entire register becomes strongly entangled.
@@ III Main Results
Here, we explore the application of GPT5.1 with an episodic memory step and no feedback. We experiment using 25 to 45 gates on 20-qubit circuits. This choice shrinks the search space, but dodges a trick recurrently applied by small models, which is the application of a high number ($\gg$ num gates) of random CNOTs. Ranging from 25 to 45 gates, we can see from Table 1 that the model improves over the initial random circuit, and can obtain up to perfect MW in some situations, as we can see for 35 gates. Nevertheless, the success probability of having an MW measure $Q(\psi)>0.8$ is $\sim 10\%$. This evidence backs up our insights, demonstrating that LLMs can sample optimal circuits beyond repeated sampling approaches, also known as the Large Language Monkeys [2]. Nevertheless, to achieve greater performances, a leap is required.
@@ IV Conclusions
Original code used for the experiment and data generated during the experiments are available at: https://github.com/AdriQD/GPT_circuits.